# Multivariate Fields of Experts

### v1

@@ IV-A Training on Denoising
$\mathcal{L}(\boldsymbol{\theta})=\frac{1}{M}\sum_{m=1}^{M}\frac{1}{\sqrt{\sigma_{m}}}\|\mathbf{x}_{m}^{*}-\mathbf{x}_{m}\|_{1},$ (18)
where $\mathbf{x}_{m}^{*}\in\operatorname{prox}_{\lambda R_{\sigma_{m}}}(\mathbf{y}_{m})$. The weighting by $1/\sqrt{\sigma_{m}}$ ensures that the model gives sufficient importance to low noise levels during training.
We follow a bilevel strategy to optimize (18). The inner optimization corresponds to the solution $\mathbf{x}^{*}_{m}$ to the proximal problem in (2), as found through accelerated gradient descent with restarts, as described in Algorithm 1. The inner loop terminates when the relative change of the iterates falls below $10^{-4}$ or after 300 iterations.
The outer optimization aims to learn $\boldsymbol{\theta}$ by minimizing the empirical loss (18). This is done using gradient-based optimization with implicit differentiation to compute the required gradients within the inner-optimization process. We use the torchdeq library [24] and apply 25 steps of the Broyden algorithm [25] to approximate the gradient of the fixed-point solution.
@@
The training is conducted on a dataset of 238,400 small patches of size $\left(40\times 40\right)$, extracted from 400 images from the BSD500 dataset [26] and scaled to $[0,1]$. This number of training images is relatively small compared to typical deep-learning-based methods, which often rely on significantly bigger datasets to achieve good generalization [27]. We use the ADAM optimizer with a batch size of 128, setting the learning rate to 0.005 for the filters $\mathbf{W}$ and the function $\boldsymbol{\mu}$, while all other parameters use a learning rate of 0.05. The model is trained for 5000 steps, and the learning rates are decayed by a factor of 0.75 every 500 steps. The full training procedure takes around 5.5 hours on a Tesla V100 GPU.
@@
We report the denoising performance of MFoE in Table I, using $K=15$ potentials of input dimension $d=4$ with filters of size $\left(11\times 11\right)$. For comparison purposes, we include TV [8], the WCRR [5] and Prox-DRUNet [16]. Note that the WCRR results differ slightly from those in the original paper, as we trained it on noise levels ranging from 0 to $50/255$, whereas the original implementation was ranging from 0 to $30/255$. We achieved this by modifying only the splines $s_{\alpha_{i}}$ from [5, Equation (17)], increasing the number of knots from 11 to 18 to maintain interval lengths as close as possible to the original configuration.
@@
As mentioned in Section III-B, we add the relaxed version of WCRR (WCRR-free) and an alternative version of our model based on the $\ell_{2}$-norm (MFoE-$\ell_{2}$). More details about their implementation can be found in Appendices -B and -C.
@@
Relaxing the 1-weak convexity constraint in WCRR does not yield performance improvements, except at the highest noise level. In contrast, the proposed MFoE model consistently outperforms the univariate baselines (WCRR and WCRR-free) as well as the multivariate variant MFoE-$\ell_{2}$. Notably, all four models employ the same number of filters $\left(Kd=60\right)$ and have roughly the same number of parameters, which ensures that the observed performance gains are not due to increased model capacity but result instead from the greater expressiveness of the proposed multivariate potentials.
@@
We present in Table II the performance of the MFoE as we increase the number of filters from $Kd=60$ to 240. Unlike the WCRR, which saturates at 60 filters, our model continues to improve and reaches a performance plateau around 180 filters. To ensure fair comparisons throughout the paper, we adopt the 60-filter variant as our default model.
@@
We further investigate the tradeoff between $K$ and $d$ in Table VII, keeping the total number of filters fixed at $Kd=60$. We observe a significant improvement when moving from $d=1$ to $d=2$, with performance peaking at $d=4$. Beyond that, performance degrades steadily, with $d=60$ performing worse than even the univariate case $d=1$.
TABLE I: Denoising performance in terms of noise level on the BSD68 test set.
$\sigma=15/255$ $\sigma=25/255$ $\sigma=50/255$ Parameters
TV 29.90 27.48 24.83 1
WCRR 31.20 28.68 25.68 $1.4\cdot 10^{4}$
WCRR-free 31.18 28.68 25.78 $1.4\cdot 10^{4}$
MFoE-$\ell_{2}$ 31.22 28.71 25.80 $1.4\cdot 10^{4}$
MFoE 31.32 28.84 25.92 $1.4\cdot 10^{4}$
Prox-DRUNet 31.70 29.18 26.13 $1.7\cdot 10^{7}$
@@
IV-B Visualization and Analysis of the Learned Model
We display the filters of the $(K=15,d=4)$ model in Figure 2. Filters within the same group tend to exhibit similar and complementary structures, while filters across different groups often vary significantly. Since the potential function depends primarily on the filter that produces the strongest response, this behavior suggests a form of invariance in the model—possibly to orientation, scale, or other transformations.
@@ IV-B Visualization and Analysis of the Learned Model
We also plot a learned potential of the $(K=30,d=2)$ model along with its two corresponding filter in Figure 4 and show two additional ones in Section -E.
Figure 4: Learned potential function and the corresponding filters.
A direct corollary of Theorem 1 is that the minimizers of our regularizer lie in the null space of $\mathbf{W}$. In Figure 5, we show the impulse response of the filter that corresponds to $\mathbf{W}^{T}\mathbf{W}$, along with the magnitude of its Fourier transform. While the impulse response is close to a Kronecker impulse, it cannot be exactly so as it must sum up to zero due to the zero-mean constraint on all filters in $\mathbf{W}$. The frequency response reveals that only the zero-frequency coefficient vanishes, which suggests that the null space of $\mathbf{W}$ is one-dimensional and consists of constant images. However, this is not entirely accurate: because our filters are zero-padded, $\mathbf{W}^{T}\mathbf{W}$ is not perfectly diagonalized by the discrete Fourier transform—although the approximation is very close, as shown in [28]. It is also intuitive that nonzero constant images are not in the null space of $\mathbf{W}$, as the discontinuities they create at the boundaries result in nonzero filter responses.
To complete the analysis, we run power iterations on $\left(\mathbf{I}-\mathbf{W}^{T}\mathbf{W}\right)$ in float64 precision to accurately estimate the smallest singular value of $\mathbf{W}$. The result, in excess of $2.5\cdot 10^{-3}$, shows that the null space of $\mathbf{W}$ is empty. Consequently, our regularizer admits a unique global minimum, the zero-valued image.
A qualitative comparison of denoising results between our model and univariate baselines reveals that our approach performs particularly well in high-frequency regions that resemble textures as shown in Figure 3. This observation aligns with techniques used in classic texture-classification frameworks [29, 30], where features are derived from the strongest response across a bank of filters with varying orientations and scales.
@@
IV-C Inverse Problems
We benchmark MFoE for image deblurring, CS-MRI, and CT. For this, we simulate measurements as
@@ IV-C Inverse Problems
where $\mathbf{H}$ is the linear operator that corresponds to the imaging modality and $\mathbf{w}\sim\mathcal{N}(\mathbf{0},\sigma_{\mathbf{w}}\mathbf{I})$.
We compare our learned regularizer with TV, the WCRR, and Prox-DRUNet. Image reconstruction with these methods requires the minimization of the objective (1). This is done via an iterative optimization algorithm which is terminated when the relative difference of the iterates is below $10^{-5}$ or when the number of iterations exceeds 1000.
@@
IV-C1 Deblurring
We evaluate the frameworks on a deblurring task, where $\mathbf{H}$ is a convolution operator. We took from [31] three kernels of size (25 $\times$ 25) shown on the top of Table III. One kernel is isotropic Gaussian and the other two are anisotropic kernels that correspond to a motion blur. We evaluate the noise levels $\sigma_{\mathbf{w}}\in\{0.01,0.03\}$. Each method is initialized with the observed blurred and noisy measurements, the hyperparameters are tuned on the Set12 dataset, and the performances are reported on the BSD68 dataset.
We report the performances in Table III. The MFoE outperforms the WCRR in every instance and manages to match the performances of Prox-DRUNet for the isotropic blur. Examples of deblurring are provided in Figure 6. There, we can see in the zoomed part that Prox-DRUNet makes a somewhat bolder prediction than the other models. This results in a higher SSIM but a slightly lower PSNR than MFoE, despite its blurrier prediction.
TABLE III: PSNR values for the deblurring on the BSD68 dataset.
@@ IV-C1 Deblurring
TV 26.74 25.62 29.16 26.17 29.36 26.05
WCRR 27.25 26.09 30.26 26.88 30.38 26.75
MFoE 27.36 26.18 30.43 27.14 30.65 27.04
Prox-DRUNet 27.34 26.17 30.55 27.24 30.78 27.15
@@
IV-C2 Compressed-Sensing MRI
We use fully sampled knee images of size $\left(320\times 320\right)$ from the fastMRI dataset [32] as ground truth, including scans with (PDFS) and without (PD) fat suppression. The corresponding measurements are obtained by a subsampling of the 2D Fourier transform ($k$-space) using a Cartesian sampling mask defined by two parameters: the acceleration factor $M_{\text{acc}}$ and the center fraction $M_{\text{cf}}$. Specifically, the central $\lfloor 320M_{\text{cf}}\rfloor$ columns of $k$-space (those corresponding to low frequencies) are fully retained, while the remaining columns are uniformly subsampled, which results in a total of $\lfloor 320/M_{\text{acc}}\rfloor$ selected columns. Gaussian noise with standard deviation $\sigma_{\mathbf{w}}=0.01$ is added to the real and the imaginary part of the measurements. Each method is initialized with the zero-filled image, the hyperparameters are tuned on a validation set of 10 images, and the performance is reported on a testing set of 50 images, all normalized to the range $[0,1]$.
The performance is reported in Table IV. Again, the MFoE outperforms the WCRR in every instance. It also manages to outperform Prox-DRUNet on the simpler instance with the PD data. Examples of reconstruction are provided in Figure 7.
TABLE IV: PSNR values for MRI reconstruction on the fastMRI dataset.
$\left(M_{\text{acc}}=4,M_{\text{cf}}=0.08\right)$ $\left(M_{\text{acc}}=8,M_{\text{cf}}=0.04\right)$

### v2

@@ IV-A Bilevel Formulation
$\min_{\boldsymbol{\theta}}\quad$ $\mathcal{L}(\boldsymbol{\theta})=\frac{1}{M}\sum_{m=1}^{M}\frac{1}{\sqrt{\sigma_{m}}}\|\mathbf{x}_{m}^{*}(\boldsymbol{\theta})-\mathbf{x}_{m}\|_{1}$ (21)
s.t. $\mathbf{x}_{m}^{*}(\boldsymbol{\theta})=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{n}}\left(\frac{1}{2}\|\mathbf{x}-\mathbf{y}_{m}\|_{2}^{2}+\lambda R_{\sigma_{m}}(\mathbf{x})\right).$ (22)
The dependence on $\sigma$ is only enforced through the parameters $\{\mu_{k}\}_{k=1}^{K}$, while the filters and other components remain the same across all noise levels. We describe their parameterization in Appendix -C.
The weighting factor $1/\sqrt{\sigma_{m}}$ ensures that the loss remains sensitive to errors at low noise levels. This is a common strategy used in the training of diffusion models [38].
IV-B Inner Optimization (Forward Pass)
The inner problem consists of the minimization of (22). Since this objective is not necessarily convex, we find an approximate stationary point $\mathbf{x}_{m}^{*}$ by running the heavy-ball with restart (see Algorithm 1) until the relative change of the iterates falls below $10^{-4}$. Since $\mathbf{\nabla}f(\mathbf{x}_{m}^{*})\approx 0$, the point $\mathbf{x}_{m}^{*}$ can be viewed as the fixed point of the gradient-step operator
$\mathbf{x}_{m}^{*}\approx\mathbf{x}_{m}^{*}-\alpha\mathbf{\nabla}f(\mathbf{x}_{m}^{*})=\mathcal{T}(\mathbf{x}_{m}^{*}).$ (23)
IV-C Outer Optimization (Backward Pass)
We need access to the gradient $\nabla_{\boldsymbol{\theta}}\mathcal{L}$ to minimize the outer objective (21) via gradient descent. Applying the chain rule, we have that
$\nabla_{\boldsymbol{\theta}}\mathcal{L}=\sum_{m=1}^{M}\left(\frac{\partial\mathbf{x}_{m}^{*}}{\partial\boldsymbol{\theta}}\right)^{T}\nabla_{\mathbf{x}_{m}^{*}}\mathcal{L}.$ (24)
As shown in the deep equilibrium framework [5], one can compute the gradient of $\mathbf{x}_{m}^{*}$ with respect to $\boldsymbol{\theta}$ using the implicit function theorem applied to (23) as
$\frac{\partial\mathbf{x}_{m}^{*}}{\partial\boldsymbol{\theta}}=\left(\mathbf{I}-\mathbf{J}_{\mathcal{T}}(\mathbf{x}_{m}^{*})\right)^{-1}\frac{\partial\mathcal{T}}{\partial\boldsymbol{\theta}},$ (25)
where $\mathbf{J}_{\mathcal{T}}$ is the Jacobian of the fixed-point operator with respect to $\mathbf{x}$.
The determination of the inverse in (25) is computationally prohibitive. Instead, we approximate it using 25 steps of the Broyden algorithm [7], as implemented in the torchdeq library [17]. This allows us to backpropagate through the equilibrium point without unrolling the entire optimization trajectory, which results in significant reduction of the memory usage.
IV-D Implementation Details
@@
The training is conducted on a dataset of 238,400 small patches of size $\left(40\times 40\right)$, extracted from 400 images from the BSD500 dataset [2] and scaled to $[0,1]$. The number of training images is relatively small compared to typical deep-learning-based methods, which often rely on significantly bigger datasets to achieve good generalization [48]. We use the ADAM optimizer with a batch size of 128 and set the learning rate to 0.005 for the filters $\mathbf{W}$ and the function $\boldsymbol{\mu}$, while all other parameters use a learning rate of 0.05. The model is trained for 5,000 steps and the learning rates are decayed by a factor of 0.75 every 500 steps. The full training procedure takes about 5 hours on a Tesla V100 GPU.
IV-E Results and Ablation Studies
We report the denoising performance of MFoE in Table I on the grayscale-converted BSD68 [25], McMaster [50] and Set14 [47] datasets. We use $K=15$ potentials of input dimension $d=4$ with filters of size $\left(11\times 11\right)$. For comparison purposes, we include TV [35], the WCRR [19], and Prox-DRUNet [21]. The TV regularizer is a classic baseline with a very simple architecture. Prox-DRUNet is the most advanced deep-learning-based regularizer. It involves three order of magnitudes more parameters than the MFoE.
@@ IV-E Results and Ablation Studies
Note that the WCRR results differ slightly from those in the original paper because we trained it on noise levels ranging from 0 to $50/255$, whereas the original implementation was ranging from 0 to $30/255$. We achieved this by modifying only the splines $s_{\alpha_{i}}$ from [19, Equation (17)], increasing the number of knots from 11 to 18 to maintain interval lengths as close as possible to the original configuration.
@@
As mentioned in Section III-B, we add the relaxed version of WCRR (WCRR-free) and an alternative version of our model based on the $\ell_{2}$-norm (MFoE-$\ell_{2}$). More details about their implementation can be found in Appendices -D and -E.
@@
The relaxation of the 1-weak convexity constraint in WCRR yields clear improvements in performance, as evidenced by WCRR-free consistently outperforming the constrained WCRR baseline. MFoE-$\ell_{2}$ achieves results comparable to the unconstrained univariate model. However, the proposed MFoE model consistently outperforms all univariate baselines (WCRR and WCRR-free) as well as the MFoE-$\ell_{2}$ variant, which confirms the benefits of the learned multivariate $\ell_{\infty}$-based potentials. Notably, all four models employ the same number of filters $\left(Kd=60\right)$ and have roughly the same number of parameters, which ensures that the observed performance gains are not due to increased model capacity but result instead from the greater expressivity of the proposed multivariate potentials. Although Prox-DRUNet maintains a performance lead, the proposed MFoE reduces this margin while offering a principled, interpretable architecture backed by formal convergence guarantees.
@@
We present in Table II the performance of the MFoE as we increase the number of filters from $Kd=60$ to 240. Unlike the WCRR, which saturates at 60 filters, our model continues to improve and reaches a performance plateau around 180 filters. To ensure fair comparisons with the WCRR throughout the paper, we adopt the 60-filter variant as our default model.
@@
We further investigate the tradeoff between $K$ and $d$ in Table VII, keeping the total number of filters fixed at the product $Kd=60$. We observe a significant improvement when moving from $d=1$ to $d=2$, with performance peaking at $d=4$. Beyond that, performance degrades steadily, with $d=60$ performing worse than even the univariate case $d=1$. We attribute the drop in performance to the nature of the $l_{\infty}$-norm. The gradient is nonzero only for the filter with the maximal response within each group. Fewer filters get updated at each step as $K$ is decreased. This dynamic can cause filters that rarely achieve the maximal response to stagnate at their initialization.
TABLE I: Denoising performance (PSNR / SSIM) on the BSD68, McMaster, and Set14 datasets.
BSD68 McMaster Set14
Method $\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$ Params.
TV 29.90 / 0.911 27.48 / 0.857 24.83 / 0.771 31.51 / 0.930 29.07 / 0.888 26.28 / 0.822 30.19 / 0.915 27.71 / 0.864 24.86 / 0.782 1
WCRR 31.20 / 0.933 28.68 / 0.890 25.68 / 0.806 33.24 / 0.952 30.61 / 0.920 27.23 / 0.849 31.74 / 0.938 29.19 / 0.900 25.93 / 0.817 $1.4\cdot 10^{4}$
WCRR-free 31.18 / 0.932 28.68 / 0.888 25.77 / 0.807 33.32 / 0.953 30.77 / 0.923 27.53 / 0.861 31.76 / 0.937 29.27 / 0.899 26.12 / 0.821 $1.4\cdot 10^{4}$
MFoE-$\ell_{2}$ 31.21 / 0.933 28.71 / 0.889 25.80 / 0.809 33.32 / 0.953 30.77 / 0.923 27.50 / 0.860 31.77 / 0.938 29.28 / 0.900 26.08 / 0.822 $1.4\cdot 10^{4}$
MFoE 31.32 / 0.935 28.84 / 0.892 25.93 / 0.812 33.53 / 0.955 31.02 / 0.926 27.73 / 0.865 31.96 / 0.940 29.54 / 0.904 26.35 / 0.827 $1.4\cdot 10^{4}$
Prox-DRUNet 31.70 / 0.940 29.18 / 0.898 26.14 / 0.815 33.92 / 0.957 31.32 / 0.927 27.97 / 0.860 32.50 / 0.945 30.06 / 0.911 26.83 / 0.835 $1.7\cdot 10^{7}$
@@
V Visualization and Analysis of the Learned Model
A qualitative comparison of our denoising results with the univariate baselines reveals that our approach performs particularly well on periodic patterns, as shown in Figure 4 on the zebra stripes.
@@ V Visualization and Analysis of the Learned Model
We attribute this improvement to the mechanics of quadrature filtering. Consider a pure harmonic signal $x(t)=\cos(\omega_{0}t)$. A single linear filter $h$ transforms this signal into $(h*x)(t)=A\cos(\omega_{0}t+\varphi)$, where $A=|H(\omega_{0})|$ and $\varphi=\arg(H(\omega_{0}))$ denote the spectral magnitude and phase, respectively. Because the response of the filter oscillates, the regularization penalty induced by a univariate potential will fluctuate spatially. This behavior may result in an uneven signal reconstruction. To avoid this effect, one may use a pair of quadrature filters $(h_{1},h_{2})$ [1]. These are built such that $|H_{1}(\omega_{0})|=|H_{2}(\omega_{0})|=A$ and $\arg(H_{2}(\omega_{0}))=\arg(H_{1}(\omega_{0}))+\frac{\pi}{2}$. By combining their output via a square root, we get
$\sqrt{(h_{1}*x)^{2}+(h_{2}*x)^{2}}$
$\quad=\sqrt{A^{2}\cos^{2}(\omega_{0}t+\varphi)+A^{2}\sin^{2}(\omega_{0}t+\varphi)}$
$\quad=A.$ (26)
This is a configuration that cannot be achieved by a univariate architecture.
Figure 3: A pair of learned filters from the $\left(K=30\text{ and }d=2\right)$ model. In the top row, a red frame highlights a location where the two filters differ most markedly. The magnitude of the Fourier transforms are shown in the bottom row.
The filters of the $(K=15,d=4)$ model displayed in Figure 2 exhibit similar and complementary structures within the same group. Those filters can work together to extract meaningful patterns in schemes that are reminiscent of quadrature-filter pairs and, more generally, steerable filterbanks [16, 41]. We show a group of two filters and their Fourier magnitudes from the $(K=30,d=2)$ model in Figure 3. Those two filters are almost identical except for a horizontal shift of a group of two pixels, as highlighted by the red frame. Because of this, they have very similar Fourier magnitudes, which suggests some quadrature mechanism within this group.
@@
We plot the learned potential of those same two filters in Figure 5. Visually, the learned potential resembles a rotated $\ell_{p}$-norm with $p<1$, a geometry characteristic of antisparsity [14]. The regularizer penalizes correlated responses less severely than independent ones. This property favors joint activations and avoids the detrimental periodic nature of individual filter responses via a coupling similar to that of quadrature filters.
Figure 5: Learned potential function and the corresponding filters. The axes correspond to the responses of the two filters.
A corollary of Theorem 1 is that the minimizers of our regularizer lie in the null space of $\mathbf{W}$. In Figure 6, we show the impulse response of the filter that corresponds to $\mathbf{W}^{T}\mathbf{W}$, along with the magnitude of its Fourier transform. While the impulse response is close to a Kronecker impulse, it cannot be exactly so as it must sum up to zero due to the zero-mean constraint on all filters in $\mathbf{W}$. The frequency response reveals that only the zero-frequency coefficient vanishes. This suggests that the null space of $\mathbf{W}$ is one-dimensional and consists of constant images.
@@
VI Inverse Problems
We benchmark MFoE for image deblurring, CS-MRI, and CT. For each task, we simulate measurements as
@@ VI Inverse Problems
where $\mathbf{H}$ is the linear operator corresponding to the imaging modality and $\mathbf{w}\sim\mathcal{N}(\mathbf{0},\sigma_{\mathbf{w}}^{2}\mathbf{I})$.
We compare our learned regularizer with TV, WCRR, and Prox-DRUNet. To reconstruct the images, we minimize the objective (1) using the solvers recommended in the respective original works; namely, proximal gradient descent for TV, safeguarded accelerated gradient descent [19] for WCRR, Douglas-Ratchford splitting [12] for Prox-DRUNet, and the heavy-ball method with restart (Algorithm 1) for MFoE. In all cases, the algorithm is terminated when the relative difference between consecutive iterates falls below $10^{-5}$ or when the number of iterations exceeds 1,000.
@@
VI-A Deblurring
We apply the frameworks to a deblurring task, where $\mathbf{H}$ is a convolution operator. We take three kernels of size (25 $\times$ 25) from [24], as shown on the top of Table III. The first kernel is isotropic Gaussian while the other two are anisotropic and model a motion blur. We consider the noise levels $\sigma_{\mathbf{w}}\in\{0.01,0.03\}$. Each method is initialized with the observed blurred and noisy measurements. The hyperparameters are tuned on the Set12 [49] dataset and the performance is reported in Table III on the BSD68 dataset.
The MFoE outperforms the WCRR in every instance and manages to match the performance of Prox-DRUNet for the isotropic blur. Examples of deblurring are provided in Figure 7. There, we can see in the zoomed part that Prox-DRUNet makes a somewhat bolder prediction than the other models. This results in a higher SSIM but a lower PSNR than MFoE, despite the blurrier prediction of MFoE.
TABLE III: Quantitative results (PSNR / SSIM) for deblurring on the BSD68 dataset.
@@ VI-A Deblurring
TV 26.74 / 0.846 25.62 / 0.803 29.16 / 0.902 26.17 / 0.820 29.36 / 0.904 26.05 / 0.815
WCRR 27.25 / 0.859 26.09 / 0.819 30.26 / 0.920 26.88 / 0.838 30.38 / 0.922 26.75 / 0.833
MFoE 27.36 / 0.860 26.18 / 0.819 30.43 / 0.921 27.14 / 0.848 30.65 / 0.926 27.04 / 0.844
Prox-DRUNet 27.34 / 0.859 26.17 / 0.820 30.55 / 0.924 27.24 / 0.852 30.78 / 0.928 27.15 / 0.848
@@
VI-B Compressed-Sensing MRI
We use fully sampled real-valued knee images of size $\left(320\times 320\right)$ from a fastMRI dataset [46] as ground truth, including scans with (PDFS) and without (PD) fat suppression. The corresponding measurements are obtained by a subsampling of the 2D Fourier transform ($k$-space). The Cartesian sampling mask is defined by two parameters: the acceleration factor $M_{\text{acc}}$; and the center fraction $M_{\text{cf}}$. Specifically, the central $\lfloor 320M_{\text{cf}}\rfloor$ columns of $k$-space (those corresponding to low frequencies) are fully retained, while the remaining columns are uniformly subsampled, which results in a total of $\lfloor 320/M_{\text{acc}}\rfloor$ selected columns. Gaussian noise with standard deviation $\sigma_{\mathbf{w}}=0.01$ is added to the real and the imaginary parts of the measurements. Each method is initialized with the zero-filled image, the hyperparameters are tuned on a validation set of 10 images, and the performance is reported on a testing set of 50 images, all normalized to the range $[0,1]$.
The performance is reported in Table IV. Again, the MFoE outperforms the WCRR in every instance. It also manages to outperform Prox-DRUNet on the $\left(M_{\text{acc}}=4,M_{\text{cd}}=0.08\right)$ instance with the PD data. Examples of reconstruction are provided in Figure 8.
TABLE IV: Quantitative results (PSNR / SSIM) for MRI reconstruction on the fastMRI dataset.
$(M_{\text{acc}}=4,M_{\text{cf}}=0.08)$ $(M_{\text{acc}}=8,M_{\text{cf}}=0.04)$